A 197g ball is tied to a string. It is pulled to an angle of 8.0? and released to swing as a pendulum. A student with a stopwatch finds that 21 oscillations take 11s .

How long is the string?
Answer #1

Since the angle pulled back is small, we can use the approximation:
T = 2??(L/g).

Solving this for L gives:
L = (T^2*g)/(4?^2).

Since the period is21/11= 1.909 s and g = 9.80 m/s^2:
L = (T^2*g)/(4?^2)
= [(1.909 s)^2(9.80 m/s^2)]/(4?^2)
= 0.9056m = 90.56 cm.